OpenStudy (anonymous):

A formula for the area of a square is A=s^2, where s is a side. Suppose the area of a square is given to the right. A. ( I have this one ) B. Find s(x)? C. Find s(3)?

OpenStudy (anonymous):

A=9x^2-12x+4

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

A=s^2=9x^2-12x+4=(3x)^2 -2*3x*2+(2)^2=(3x-2)^2

OpenStudy (anonymous):

So, s^2=(3x-2)^2 s=3x-2 s(x)=3x-2

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

yep what about s(3)

OpenStudy (anonymous):

would it be 7?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

ok i have another one but it's hard to describe on here.. what do u suggest i do?

OpenStudy (anonymous):

from which topic?

OpenStudy (anonymous):

it's math but it's a word problem

OpenStudy (anonymous):

Just copy and paste it

OpenStudy (anonymous):

i cant copy it.. it's on a program called math xl and it want let me copy and paste

OpenStudy (anonymous):

do u want me to type it right here and see what u think... the problem has 3 parts to it

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

u post in another question

OpenStudy (anonymous):

and tag me

OpenStudy (anonymous):

ok give me just a sec and it will be up